# Myopically Verifiable Probabilistic Certificates for Safe Control and Learning

### v1

@@ IV-B2 Forward Convergence
[…]
$\Theta_{x}(T_{t})$ $:=\sup\{\phi(X_{\tau})\in\mathbb{R}:\tau\in[0,T_{t}],X_{0}=x\},$ (17)
$\Psi_{x}(L_{t})$ $:=\inf\{\tau\in\mathbb{R}_{+}:\phi(X_{\tau})\geq L_{t},X_{0}=x\}.$ (18)
Here, $\Theta_{x}(T_{t})$ indicates the distance to the boundary of the safe set $\partial{\mathcal{C}}(0)$, and $\Psi_{x}(L_{t})$ is the duration for the state to enter the set ${\mathcal{C}}(L_{t})$ for the first time. We can also rewrite (16) using the two random variables (17) and (18) as
$=\>\mathbb{P}\left(\ \exists\tau\in[0,T_{t}]\text{ s.t. }X_{\tau}\in{\mathcal{C}}(L_{t})\ |\ X_{0}=x\ \right)$ (20)
$=\>\mathbb{P}(\Theta_{x}(T_{t})\geq L_{t})$ (21)
$=\>\mathbb{P}(\Psi_{x}(L_{t})\leq T_{t}).$ (22)
@@
$P_{3}(x,L,T,\phi):=\mathbb{P}\{X_{\tau}\in{\mathcal{C}}(L)$ (23)
[…]
@@
$P_{4}(x,L,T,\phi):=\sup_{N\in\mathcal{N}}\mathbb{P}\{X_{\tau}\in{\mathcal{C}}(L)$ (24)
@@
IV-C Design Goals
This objective of this paper is to ensure either long-term safety/avoidance or finite-time eventuality/reachability. The objective is mathematically defined as follows. For simplicity of notations, we define the following probability
$P(x,L,T,\phi):=\begin{cases}P_{1}(x,L,T,\phi),\quad\text{for type 1}\\ P_{2}(x,L,T,\phi),\quad\text{for type 2}\\ P_{3}(x,L,T,\phi),\quad\text{for type 3}\\ P_{4}(x,L,T,\phi),\quad\text{for type 4}\end{cases}$ (25)
@@ IV-C Design Goals
where $\epsilon\in(0,1)$ is the pre-specified risk tolerance. The expectation (26) is taken over the distribution of $X_{t}$ and the future trajectories $\{X_{\tau}\}_{\tau\in(t,t+T_{t}]}$ conditioned on the initial distribution of $X_{0}$. The distribution of $X_{t}$ is generated based on the closed-loop system of (2) and (4), whereas the distribution of $\{X_{\tau}\}_{\tau\in(t,t+T_{t}]}$ are allowed to be defined in two different ways based on the design choice: the closed-loop system of (2) and (3) or the closed-loop system of (2) and (4).
@@
We consider either fixed time horizon problem or receding time horizon problem. In the fixed time horizon problem, safety is evaluated at each time $t$ for a time interval $[t,t+H]$ of fixed length. In the receding time horizon problem, we evaluate, at each time $t$, safety only for the remaining time $[t,H]$ given a fixed horizon. The outlook time horizon for each case is given by
$T_{t}$ =$=\begin{cases}H-t,&\text{for fixed time horizon,}\\ H,&\text{for receding time horizon.}\end{cases}$ (28)
@@
V Probabilistic Safety Certificate
Here, we present a sufficient condition to achieve the safety requirements in subsection V-A. Then we provide efficient computation methods to calculate the safety condition in subsection V-B. Finally, we propose two safe control strategies and conclude the overall framework in subsection V-C.
@@ V Probabilistic Safety Certificate
$\tilde{f}(Z_{t})$ $:=\begin{bmatrix}f_{T}\\ f_{\ell}(L_{t})\\ f_{\phi}(X_{t})\\ f(X_{t})\end{bmatrix}\in\mathbb{R}^{(n+3)},$ (32)
$\tilde{g}(Z_{t})$ $:=\begin{bmatrix}\mathbf{0}_{2\times m}\\ {\mathcal{L}}_{g}\phi(X_{t})\\ g(X_{t})\end{bmatrix}\in\mathbb{R}^{(n+3)\times m},$ (33)
$\tilde{\sigma}(Z_{t})$ $:=\begin{bmatrix}\mathbf{0}_{2\times\xi}\\ {\mathcal{L}}_{\sigma}\phi(X_{t})\\ \sigma(X_{t})\end{bmatrix}\in\mathbb{R}^{(n+3)\times\xi}.$ (34)
@@
$f_{T}$ :=$:=\begin{cases}0,&\text{in fixed time horizon,}\\ -1,&\text{in receding time horizon,}\end{cases}$ (35)
@@
$f_{\phi}(X_{t})$ $:={\mathcal{L}}_{f}\phi(X_{t})+\frac{1}{2}\text{tr}\big{(}\left[\sigma(X_{t})\right]\left[\sigma(X_{t})\right]^{\intercal}\operatorname{Hess}{\phi}(X_{t})\big{)}.$ (36)
@@
V-A Conditions to Assure Safety
We consider the following probabilistic quantity: [^4] Recall from Section IV-B that whenever we take the probabilities (and expectations) over paths, we assume that the probabilities are conditioned on the initial condition $X_{0}=x$.
@@ V-A Conditions to Assure Safety
where the probability is taken over the same distributions of $\{X_{\tau}\}_{\tau\in[t,T_{t}]}$ that are used in the safety requirement (26). The values of $T_{t}$ and $L_{t}$ (known and deterministic) are defined in 28 and 29 depending on the design choice of receding/fixed time-horizon and fixed/varying margin.
Additionally, we define the mapping $D_{\mathbf{F}}:\mathbb{R}^{n+3}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$ as [^5] See Remark 1 for the notation for Lie derivative.
$\begin{split}D_{\mathbf{F}}(Z_{t},U_{t}):=&\ {\mathcal{L}}_{\tilde{f}}\mathbf{F}(Z_{t})+\left({\mathcal{L}}_{\tilde{g}}\mathbf{F}(Z_{t})\right)U_{t}\\ &+\frac{1}{2}\text{tr}\left(\left[\tilde{\sigma}(Z_{t})\right]\left[\tilde{\sigma}(Z_{t})\right]^{\intercal}\operatorname{Hess}{\mathbf{F}}(Z_{t})\right).\end{split}$ (38)
From Itô’s Lemma, [^6] Itô’s Lemma is stated as below: Given a $n$-dimensional real valued diffusion process $dX=\mu dt+\sigma dW$ and any twice differentiable scalar function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, one has $df=\left({\mathcal{L}}_{\mu}f+\frac{1}{2}\operatorname{tr}\left(\sigma\sigma^{\intercal}\operatorname{Hess}{f}\right)\right)dt+{\mathcal{L}}_{\mu}\sigma dW.$ the mapping (38) essentially evaluates the value of the infinitesimal generator of the stochastic process $Z_{t}$ acting on $\mathbf{F}$: i.e., $A\mathbf{F}(Z_{t})=D_{\mathbf{F}}(Z_{t},U_{t})$ when the control action $U_{t}$ is used.
@@
[…]
$\mathbb{E}\left[\mathbf{F}(Z_{t})\right]\geq 1-\epsilon$ (40)
for all time $t\in\mathbb{R}_{+}$.
Proof (theorem 1).
@@
$D_{0}$ =$=\left\{\mathbf{F}(Z_{\tau})<1-\epsilon\right\},$ (43)
$D_{1}$ =$=\left\{\mathbf{F}(Z_{\tau})\geq 1-\epsilon\right\},$ (44)
@@
$q_{0}$ $=\mathbb{P}(D_{0}),$ (47)
$q_{1}$ $=\mathbb{P}(D_{1}).$ (48)
The left hand side of 41 can then be written as
@@
[…]
@@
[…]
Here, (57) is due to (47) and (48); (58) is obtained from Jensen’s inequality [45] for concave function $\alpha$; (59) is based on (45) and (46); (60) is given by the assumptions on function $\alpha$; and (61) is due to (55). Thus, we showed that (41) implies (42).
@@
Let $y\colon\mathbb{R}_{+}\to\mathbb{R}$ be a real-valued differentiable function that satisfies
$\frac{d}{dt}y_{t}\geq 0\quad\text{whenever $y_{t}\leq L$}.$ (66)
Additionally, we assume $y_{0}>L$. Then
$y_{t}\geq L\quad\text{for all $t\in\mathbb{R}_{+}$.}$ (67)
Proof (lemma 1).
Suppose there exists $b\in\mathbb{R}_{+}$ such that $y_{b}<L$. By the intermediate value theorem, there exists $a\in(0,b)$ such that $y_{a}=L$, and $y_{t}<L$ for all $t\in(a,b]$. Next, by the mean value theorem, there exists $\tau\in(a,b)$ such that $(dy_{t}/dt)|_{t=\tau}=(y_{b}-y_{a})/(b-a)<0$. This contradicts condition (66). ∎
V-B Efficient Computation of Long-term Probabilities
In this section, we provide computation methods to calculate the probability of the safety-related events. While the proposed safe control technique works for both stochastic and deterministic control policies, we here present computation methods for the cases of deterministic policies which are more commonly considered. Before introducing computation methods for controller implementation, we first consider a discretized version of dynamical system (2) for sampling time $\Delta t$:
@@ V-B Efficient Computation of Long-term Probabilities
V-B1 Long-term safety and finite time eventuality
Importance sampling techniques provide us with efficient computation methods to calculate the probabilities when the nominal control is fixed [46, 47]. [^9] For simplicity, we assume $g(X)=\sigma(X)$ as in the path integral formulation. Assume $N\in\mathcal{N}_{d}$ is the nominal controller that we are interested in, and $N_{s}\in\mathcal{N}_{d}$ is another controller that we have sampled data with. Let $\mathcal{P}_{N}$ denote the measure over the classical Wiener space $\Omega=C([0,T];\mathbb{R}^{n})$ induced by the dynamics (68) with $U_{k}=N(X_{k})$, and $\mathcal{P}_{N_{s}}$ denote the measure associated with the process with $U_{k}=N_{s}(X_{k})$. Note that $N_{s}$ can output zero values where the process is uncontrolled. Let $\tilde{U}_{k}=N(X_{k})-N_{s}(X_{k})$ denote the difference between the nominal controller and the sampled controller, where $X_{k}$ is the state of the closed-loop system with $N_{s}(X_{k})$.
Let $S_{l}=\mathds{1}(X_{k}\in\mathcal{C}(L),\forall k\in\{1,2,\cdots,T_{d}\})$ and $S_{f}=\mathds{1}(X_{k}\in\mathcal{C}(L),\text{ for some }k\in\{1,2,\cdots,T_{d}\})$ be the indicator function of safety events for long-term safety and finite-time eventuality. We use $S$ to denote $S_{l}$ or $S_{f}$ given context. Through importance sampling one can sample $\mathbf{F}(z)$ based on controller $N_{s}$ and take weighted expectation with coefficients $\frac{d\mathcal{P}_{N}}{d\mathcal{P}_{N_{s}}}$ to estimate $\mathbf{F}(z)$ for $N$ as follows [^10] Recall that $Z_{t}=[T,L,\phi(X_{t}),X_{t}]$ as defined in (30). Here, we assume that $T$ and $L$ are fixed. The lowercase letters $x$, $z$ are used to denote the specific realizations of $X_{t}$ and $Z_{t}$.
$\mathbf{F}(z):=\mathbb{P}\left(S\mid X_{0}=x\right)$ (69)
@@ V-B1 Long-term safety and finite time eventuality
With (69) and (70), one can estimate the safety probability for any controller $N$ by sampling controller $N_{s}$. Note that with samples on $N_{s}$, we can estimate $\mathbf{F}(z)$ for arbitrary $N$ using (69). Plus, when $N$ and $N_{s}$ are both represented by neural networks, we can still calculate $\frac{d\mathcal{P}_{N}}{d\mathcal{P}_{N_{s}}}$ by sampling $\tilde{U}_{k}$ at each time step. Specifically, to do the calculation, one can sample the process with controller $N_{s}$ for $n_{\text{sample}}$ times. We use
$s_{i}:=\mathds{1}(X_{k}\in\mathcal{C}(L),\forall k\in\{1,2,\cdots,T_{d}\})$ (71)
@@
$s_{i}:=\mathds{1}(X_{k}\in\mathcal{C}(L),\text{ for some }k\in\{1,2,\cdots,T_{d}\})$ (72)
for finite-time eventuality. This way we can estimate (69) with
@@
Input: $N$, $N_{s}$, $T_{d}$, $n_{\text{sample}}$, $z=[T,L,\phi(x),x]$
for $i\in\{1,\cdots,n_{\text{sample}}\}$ do
@@
         Calculate $X_{k+1}$ using the system dynamics (68) with $U_{k}$
@@
Physics-informed learning can also be used to further improve the efficiency of computation for safety-related probabilities in a way that generalizes to unsampled states or time horizons [49, 50, 51].
V-B2 Long-term avoidance and finite-time reachability

### v2

@@ 4.2.2 Forward Convergence
for some time interval $t,t+T_{t}]$ conditioned on an initial condition $x\in{\mathcal{C}}(L_{t})^{c}$.
@@
$P_{3}(x,L,T,\phi)=\mathbb{P}\{X_{\tau}\in{\mathcal{C}}(L)$ (11)
 $\quad\text{ for some }\tau\in[t,t+T]\mid X_{t}=x\},$
@@
$P_{4}(x,L,T,\phi)=\sup_{N\in\mathcal{N}}\mathbb{P}\{X_{\tau}\in{\mathcal{C}}(L)$ (12)
 $\quad\text{ for some }\tau\in[t,t+T]\mid X_{t}=x\},$
Note that one can define the safety-related probabilities with regard to multiple safety requirements with different horizons of interest. Such generalization is straightforward thus is not explicitly discussed in the paper due to space limit.
4.3 Design Goals
The objective of this paper is to ensure either long-term safety/avoidance or finite-time eventuality/reachability. The objective is mathematically defined as follows. For simplicity of notations, we define the following probability
@@ 4.3 Design Goals
where $\epsilon\in(0,1)$ is the pre-specified risk tolerance. The expectation (14) is taken over the distribution of $X_{t}$ and the future trajectories $\{X_{\tau}\}_{\tau\in(t,t+T_{t}]}$ conditioned on the initial distribution of $X_{0}$. The distribution of $X_{t}$ is generated based on the closed-loop system of (2) and (4), whereas the distribution of $\{X_{\tau}\}_{\tau\in(t,t+T_{t}]}$ is allowed to be defined in two different ways based on the design choice: the closed-loop system of (2) and (3) or the closed-loop system of (2) and (4).
@@
We consider either fixed time horizon problem or receding time horizon problem, following the definition in  [55]. In the receding time horizon problem, safety is evaluated at each time $t$ for a time interval $[t,t+H]$. In the fixed time horizon problem, we evaluate, at each time $t$, safety only for the remaining time $[t,H]$. The outlook time horizon for each case is given by
$T_{t}$ $=\begin{cases}H,&\text{for receding time horizon,}\\ H-t,&\text{for fixed time horizon.}\end{cases}$ (16)
@@
5 Probabilistic Safety Certificate
Here, we present a sufficient condition to achieve the safety requirements in subsection 5.1. Then we provide efficient computation methods to calculate the safety condition in subsection 5.2. Finally, we propose two safe control strategies and conclude the overall framework in subsection 5.3.
@@ 5 Probabilistic Safety Certificate
$\tilde{f}(Z_{t})$ $:=[f_{T},f_{\ell}(L_{t}),f_{\phi}(X_{t}),f(X_{t})^{\top}]^{\top}\in\mathbb{R}^{(n+3)},$ (20)
$\tilde{g}(Z_{t})$ $:=[\mathbf{0}_{m\times 2},({\mathcal{L}}_{g}\phi(X_{t}))^{\top},g(X_{t})^{\top}]^{\top}\in\mathbb{R}^{(n+3)\times m},$ (21)
$\tilde{\sigma}(Z_{t})$ $:=[\mathbf{0}_{\xi\times 2},({\mathcal{L}}_{\sigma}\phi(X_{t}))^{\top},\sigma(X_{t})^{\top}]^{\top}\in\mathbb{R}^{(n+3)\times\xi}.$ (22)
@@
$f_{T}$ $:=\begin{cases}0,&\text{in receding time horizon,}\\ -1,&\text{in fixed time horizon,}\end{cases}$ (23)
@@
$f_{\phi}(X_{t})$ $:={\mathcal{L}}_{f}\phi(X_{t})+\frac{1}{2}\text{tr}\big(\left[\sigma(X_{t})\right]\left[\sigma(X_{t})\right]^{\top}\operatorname{Hess}{\phi}(X_{t})\big).$ (24)
@@
5.1 Conditions to Assure Safety
We consider the following probabilistic quantity: [^3] Recall from Section 4.2 that whenever we take the probabilities (and expectations) over paths, we assume that the probabilities are conditioned on the initial condition $X_{0}=x$.
@@ 5.1 Conditions to Assure Safety
where the probability is taken over the same distributions of $\{X_{\tau}\}_{\tau\in[t,T_{t}]}$ that are used in the safety requirement (14). The values of $T_{t}$ and $L_{t}$ (known and deterministic) are defined in ˜16 and 17 depending on the design choice of receding/fixed time-horizon and fixed/varying margin. Additionally, we define the mapping $D_{\mathbf{F}}:\mathbb{R}^{n+3}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$ as [^4] See Remark 1 for the notation for Lie derivative.
$\begin{split}D_{\mathbf{F}}(Z_{t},U_{t}):=&\ {\mathcal{L}}_{\tilde{f}}\mathbf{F}(Z_{t})+\left({\mathcal{L}}_{\tilde{g}}\mathbf{F}(Z_{t})\right)U_{t}\\ &+\frac{1}{2}\text{tr}\left(\left[\tilde{\sigma}(Z_{t})\right]\left[\tilde{\sigma}(Z_{t})\right]^{\top}\operatorname{Hess}{\mathbf{F}}(Z_{t})\right).\end{split}$ (26)
From Itô’s Lemma, [^5] Itô’s Lemma is stated as below: Given a $n$-dimensional real valued diffusion process $dX=\mu dt+\sigma dW$ and any twice differentiable scalar function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, one has $df=\left({\mathcal{L}}_{\mu}f+\frac{1}{2}\operatorname{tr}\left(\sigma\sigma^{\top}\operatorname{Hess}{f}\right)\right)dt+{\mathcal{L}}_{\mu}\sigma dW.$ the mapping (26) essentially evaluates the value of the infinitesimal generator of the stochastic process $Z_{t}$ acting on $\mathbf{F}$: i.e., $A\mathbf{F}(Z_{t})=D_{\mathbf{F}}(Z_{t},U_{t})$ when the control action $U_{t}$ is used.
@@
Consider the closed-loop system of (2) and (4). If system (2) originates at $X_{0}=x$ with $\mathbf{F}(z)>1-\epsilon$, and the control action satisfies (27) at all time, then the following condition holds for all time $t\in\mathbb{R}_{+}$: [^6] Here, the expectation is taken over $X_{t}$ conditioned on $X_{0}=x$, and $\mathbf{F}$ in 25 gives the probability of forward invariance/convergence of the future trajectories $\{X_{\tau}\}_{(t,t+T_{t}]}$ starting at $X_{t}$.
$\mathbb{E}\left[\mathbf{F}(Z_{t})\right]\geq 1-\epsilon.$ (28)
Proof (theorem˜1).
@@
$D_{0}$ $=\left\{\mathbf{F}(Z_{\tau})<1-\epsilon\right\},\quad D_{1}=\left\{\mathbf{F}(Z_{\tau})\geq 1-\epsilon\right\},$ (31)
@@
$q_{0}$ $=\mathbb{P}(D_{0}),\quad q_{1}=\mathbb{P}(D_{1}).$ (34)
The left hand side of ˜29 can then be written as
@@
$\mathbb{E}\left[\mathbf{F}(Z_{\tau})\mid D_{0}\right]$ $<1-\epsilon\ \text{and}\ \mathbb{E}\left[\mathbf{F}(Z_{\tau})\mid D_{1}\right]$ $\geq 1-\epsilon,$ (36)
@@
$\ \ \ \ \ \leq\ 0.$ (47)
Here, (43) is due to (34); (44) is obtained from Jensen’s inequality [56] for concave function $\alpha$; (45) is based on (32) and (33); (46) is given by the assumptions on function $\alpha$; and (47) is due to (41). Thus, we showed that (29) implies (30).
@@
Let $y\colon\mathbb{R}_{+}\to\mathbb{R}$ be a real-valued differentiable function that satisfies ${d\over dt}y_{t}\geq 0$ whenever $y_{t}\leq L$. If $y_{0}>L$, then $y_{t}\geq L$ for all $t\in\mathbb{R}_{+}$.
Proof (lemma˜1).
Suppose there exists $b\in\mathbb{R}_{+}$ such that $y_{b}<L$. By the intermediate value theorem, there exists $a\in(0,b)$ such that $y_{a}=L$, and $y_{t}<L$ for all $t\in(a,b]$. Next, by the mean value theorem, there exists $\tau\in(a,b)$ such that $(dy_{t}/dt)|_{t=\tau}=(y_{b}-y_{a})/(b-a)<0$. This contradicts condition that ${d\over dt}y_{t}\geq 0$ whenever $y_{t}\leq L$. ∎
Theorem 1 says that with the proposed safety certificate (27), long-term safety can be guaranteed with high probability $1-\epsilon$ at all times.
5.2 Efficient Computation of Long-term Probabilities
In this section, we provide computation methods to calculate the probability of the safety-related events. While the proposed safe control technique works for both stochastic and deterministic control policies, we present here computation methods for the more commonly considered cases of deterministic policies. Before introducing computation methods for controller implementation, we first consider a discretized version of dynamical system (2) for sampling time $\Delta t$:
@@ 5.2 Efficient Computation of Long-term Probabilities
5.2.1 Long-term safety and finite time eventuality
Importance sampling techniques provide efficient computation methods to calculate the probabilities when the nominal control is fixed [57, 58]. [^8] For simplicity, we assume $g(X)=\sigma(X)$ as in the path integral formulation. Assume $N\in\mathcal{N}_{d}$ is the nominal controller of interest, and $N_{s}\in\mathcal{N}_{d}$ is another controller that we have sampled data with. Let $\mathcal{P}_{N}$ denote the measure over the classical Wiener space $\Omega=C([0,T];\mathbb{R}^{n})$ induced by the dynamics (52) with $U_{k}=N(X_{k})$, and $\mathcal{P}_{N_{s}}$ denote the measure associated with the process with $U_{k}=N_{s}(X_{k})$. Note that $N_{s}$ can output zero values where the process is uncontrolled. Let $\tilde{U}_{k}=N(X_{k})-N_{s}(X_{k})$ denote the difference between the nominal controller and the sampled controller, where $X_{k}$ is the state of the closed-loop system with $N_{s}(X_{k})$.
Let $S_{l}=\mathds{1}(X_{k}\in\mathcal{C}(L),\forall k\in\{1,2,3,\cdots,T_{d}\})$ and $S_{f}=\mathds{1}(X_{k}\in\mathcal{C}(L),\text{ for some }k\in\{1,2,3,\cdots,T_{d}\})$ be the indicator function of safety events for long-term safety and finite-time eventuality. We use $S$ to denote $S_{l}$ or $S_{f}$ given context. Through importance sampling, one can sample $\mathbf{F}(z)$ based on controller $N_{s}$ and take weighted expectation with coefficients $\frac{d\mathcal{P}_{N}}{d\mathcal{P}_{N_{s}}}$ to estimate $\mathbf{F}(z)$ for $N$ as follows [^9] Recall that $Z_{t}=[T,L,\phi(X_{t}),X_{t}]$ as defined in (18). Here, we assume that $T$ and $L$ are fixed. The lowercase letters $x$, $z$ are used to denote the specific realizations of $X_{t}$ and $Z_{t}$.
$\mathbf{F}(z)=\mathbb{P}\left(S\mid X_{0}=x\right)$ (53)
@@ 5.2.1 Long-term safety and finite time eventuality
With (53) and (54), one can estimate the safety probability for any controller $N$ by sampling controller $N_{s}$. Note that with samples on $N_{s}$, we can estimate $\mathbf{F}(z)$ for arbitrary $N$ using (53). Plus, when $N$ and $N_{s}$ are both represented by neural networks, we can still calculate $\frac{d\mathcal{P}_{N}}{d\mathcal{P}_{N_{s}}}$ by sampling $\tilde{U}_{k}$ at each time step. Specifically, to perform the calculation, one can sample the process with controller $N_{s}$ for $n_{\text{sample}}$ times. We use
$s_{i}:=\mathds{1}(X_{k}\in\mathcal{C}(L),\forall k\in\{1,2,3,\cdots,T_{d}\})$ (55)
@@
$s_{i}:=\mathds{1}(X_{k}\in\mathcal{C}(L),\text{ for some }k\in\{1,2,3,\cdots,T_{d}\})$ (56)
for finite-time eventuality. In this way, we can estimate (53) with
@@
Input: NN, $N_{s}$, $T_{d}$, $n_{\text{sample}}$, $z=[T,L,\phi(x),x]$
for $i\in\{1,2,3,\cdots,n_{\text{sample}}\}$ do
@@
   Calculate $X_{k+1}$ using dynamics (52) with $U_{k}$
@@
Physics-informed learning can also be used to further improve the efficiency of computation for safety-related probabilities in a way that generalizes to unsampled states or time horizons [59, 60, 61, 62].
5.2.2 Long-term avoidance and finite-time reachability